OpenStudy (anonymous):

Help Please!!! Rewrite with rational exponents. (^9√9x^8y)^7

OpenStudy (mertsj):

\[\sqrt[9]9x ^{8y ^{7}}{?}\]

OpenStudy (mertsj):

That surely isn't what you meant. Could you be more clear?

OpenStudy (anonymous):

I'm trying to use the equation tab

OpenStudy (anonymous):

\[\left( \sqrt[9]{9x ^{8}y} \right)^{7}\]

OpenStudy (mertsj):

Ok. Just remember that in a rational (fractional) exponent, the top number is the power and the bottom number is the root. So lets just deal with the inside part to begin with.

OpenStudy (anonymous):

ok

OpenStudy (mertsj):